BLANK would like to organize JHO as either an LLC (taxed as a sole proprietorship) or a C corporation. In either form, the entity is expected to generate an 10 percent annual before-tax return on a $590,000 investment. BLANK'S marginal income tax rate is 35 percent and her tax rate on dividends and capital gains is 15 percent. BLANK will also pay a 3.8 percent net investment income tax on dividends and capital gains she recognizes. If BLANK organizes JHO as an LLC, BLANK will be required to pay an additional 2.9 percent for self-employment tax and an additional 0.9 percent for the additional Medicare tax. Further, she is eligible to claim the full deduction for qualified business income. Assume that JHO will pay out all of its after-tax earnings every year as a dividend if it is formed as a C corporation. (Do not round intermediate calculations and round your final answers to the nearest whole dollar.)

How much cash after taxes would BLANK receive from her investment in the first year if JHO is organized as either an LLC or a C corporation?

LLC= ?

C CORP= $37,847

What is the overall tax rate on JHO’s income in the first year if JHO is organized as an LLC or as a C corporation? (Round your final answers to 2 decimal places.)

LLC= ?

C CORP= 35.86%
Answer #1

Part A

LLC

$40686

C Corp.

$37847

Part B

LLC

31.04%

C Corp.

35.85%

LLC

Description

C corp.

Description

1.

Pretax earnings

59000

(590000*10%)

59000

(590000*10%)

2.

Entity level tax rate

0%

21%

3.

Entity level tax

0

12390

(1) × (2)

4.

Earnings after-entity-level tax

59000

(1) – (3)

46610

(1) – (3)

5.

QBI deduction

(11800)

(4) x 20%

NA

6.

Deduction for 50 percent of SE tax

(790)

(1) × 0.9235 ×0.029 × 0.5

NA

7.

Net income taxable to owner

46410

(4) + (5) + (6)

46610

(4) distributed as dividend

8.

Income tax paid by owner

16244

(7) × 0.35

6992

(7) × 0.15

9.

Self-employment tax

1580

(1) × 0.9235 × 0.029

NA

10.

Additional Medicare tax/Net investment income tax

490

(1) × 0.9235 × 0.009 additional Medicare tax

1771

(7) × 0.038 net investment income tax

11.

Owner-level tax

18314

(8) + (9) + (10)

8763

(8) + (10)

After-tax cash flow

$40686

(1) – (11)

$37847

(7) – (11)

Part B

11.

Owner-level tax

18314

(8) + (9) + (10)

8763

(8) + (10)

After-tax cash flow

$40686

(1) – (11)

$37847

(7) – (11)

Overall tax rate

31.04%

(11)/(1)

35.85%

[(3) + (11)]/(1)

LLC = 18314/59000 = 31.04%

C corp. = (12390+8763)/59000 = 35.85%